# Towards a Robust Transport Network With Self-adaptive Network Digital Twin

### v1

@@ Abstract
[…]
@@ 1 Introduction
NDT is a platform designed to help the management and automation of physical network communication tasks based on interactions with its virtual counterpart, also called virtual twin (VTwin), at a certain periodicity and fidelity [1, 2]. This relation allows the network operator to emulate different scenarios in network planning to perform a what-if analysis [3, 4]. This analysis aims to evaluate the impact of management changes in the network infrastructure when these tests are impossible due to logistic issues or costs that make this impossible, such as in production environments [5, 2]. For that reason, this platform is considered an enabling technology to optimize 5th and 6th generation (5G/6G) network communication [2], to provide improved solutions to problems such as slicing management [6, 7], routing optimization [8, 9], resource power allocation [3], and admission control [10].
@@
Using the NDT platform in a production environment can create several challenges, particularly concerning the long-term accuracy of the digital twin (DT). It is essential to account for variability in incoming data, ensuring that the VTwin continually adapts to reflect real-world changes. When we consider a VTwin neural network model to achieve this reliability, such as a graph neural network (GNN[11], it is fundamental to have a well-defined lifecycle management (LCM) that organizes all phases, such as the planning, training, and operational of the NDT platform [12, 13, 14]. When a deployed NDT is used during the operational phase, an important research question arises concerning synchronization between the VTwin and the physical twin (PTwin). Synchronization issues are widely explored in the general context of DTs, considering applications at factory workstations, industrial Internet of Things, power systems, etc. [15, 16]. However, these questions are rarely explored considering the specific scenarios of a transport network NDT based on GNN. This context presents opportunities to explore particular challenges related to synchronization between twins, such as methods to achieve this and their synchronization patterns [15].
[…]
@@
    A novel NDT implementation based on a GNN architecture, incorporating LCM principles during the operational phase, such as the re-deployment of an updated VTwin model when necessary.
@@
    As part of our architecture, a concept drift detection module is integrated to monitor incoming network traffic features. This NDT module enables to analysis of the traffic behavior to make informed decisions about when to update the VTwin model.
@@
    Relevant NDT experiments that account for environments with varying traffic conditions, including changes in traffic patterns and levels of congestion.
  • 4.
    A use case application for network management related to SLA violation prediction comparing the performance of the proposed application with and without NDT synchronization.
The remainder of this article is organized as follows: in Section 2 we compare our work with different proposed data-driven NDT in the literature, considering various levels of implementation in the transport network domain and their characteristics related to NDT supported modules, such as the synchronization one. In Section 3, we present the proposed enhanced NDT architecture, taking into account the main functional requirements of NDT, showing the essential building blocks of each, such as the VTwin architecture, the synchronization methods based on concept drift detection techniques, and complementary modules. In Section 4, we detail the experiments conducted to evaluate the proposed methods for twin synchronization and present the key evaluation results across various topologies and traffic patterns, and a use case with SLA monitoring. Finally, Section 5 summarizes the main contributions of this work and outlines directions for future research.
@@ 2 Related Work
Table 1: Comparison between NDT characteristics of the related works and our proposed architecture with support to concept drift.
@@
Rusek et al., 2020 [8]
Happ, et al., 2021 [19]
Wang et al., 2022 [6]
Güemes-Palau et al., 2022 [20]
Ferriol-Galmés et al., 2022 [21]
Saravanan et al., 2022 [22]
Lai et al., 2023 [23]
Farreras et al., 2023 [24]
Messaoudi et al., 2023 [10]
Shin et al., 2023 [25]
Yu et al., 2023 [26]
Ferriol-Galmés et al., 2023 [11]
Zalat et al., 2024 [27]
Modesto et al., 2024 [28]
Farreras et al., 2025 [7]
Aben-Athar et al., 2025 [4]
@@
Recent studies point DTs to an important technology for the 5G/6G networks [29, 30, 31, 32, 33]. This aligns with their need to establish reliable environments for developing and testing their critical parts, such as wired environments [10], where DTs can provide excellent support. However, during this wave of adoption, the intersection of DT and networking—still an emerging area of study—led to a diverse range of approaches. Many of these did not align fully with the definitions of an NDT as systematized by standards organizations [34, 35] and industry stakeholders [2, 36]. Consequently, related work found in the literature presents limited analyses of practical aspects. For example, regarding the system lifecycle or the twinning rates, the latter refers to the rate of synchronization between the PTwin current state and its corresponding representation in the VTwin [37].
More specifically, much of the literature regarding NDTs of the transport domain focused on using machine learning (ML)/surrogate model-based methods to generate the VTwin. The increase in popularity of this DT approach, which can also be called data-driven [12], can be traced back to the work of Rusek et al. [8], which established the usage of a GNN model called RouteNet for predicting path-level metrics in routing optimization and network upgrade use cases. The usage of this model was further investigated by several works, which were built on the previously established RouteNet [19, 22, 24, 38, 11, 10, 28].
Recently, the work of Ferriol-Galmés et al. [11] utilized a modified version of RouteNet, called RouteNet-Fermi, to model network communication with various topologies, traffic distributions, and queuing configurations, enabling the prediction of packet loss, delay, and jitter. In this work, since all data is provided from a packet-level simulator, there are no two-way interactions between PTwin and VTwin. Therefore, the case presented in the article does not constitute a closed-loop NDT. Moreover, the investigations did not address the possible performance degradation of the VTwin in cases of data and/or concept drifts [39]. In this context, the cases presented in this work and several others built on RouteNet [19, 22, 24, 38, 11] did not address the robustness aspect of an NDT, as they did not explore its lifecycle either.
The exception to this lack of closed-loop implementation in works based on RouteNet is the works in Messaoudi et al. [10], Zalat et al. [27], and Aben-Athar et al. [4], which are, to the best of our knowledge, the first closed-loop environments considering physical layers interacting with an artificial intelligence (AI) model. The works in [10, 4] feed an application module tailored to operate over the physical counterpart, defined by virtual switches on Mininet [40]. The work in  [27] took the same approach, but considered a small-scale physical testbed. However, a noteworthy aspect of this work is that there are elements for twin synchronization, which are based on a trigger-based online learning approach. For the other two works, despite closed-loop interactions between the physical and virtual domains, there is no DT lifecycle management controlling the rate at which both parts synchronize and update their respective states, for instance, through a retraining process. In this context, the authors in [10, 4] ran the traffic only once, and after completing it, they performed no further operations to detect whether the VTwin performance would remain acceptable.
In addition to these, other works in the literature either proposed purely GNN-based approaches, such as FlowDT [21] or GNNetSlice [7], or adopted different ML models as their VTwins. For example, Lai et al. [23] adopted a model named eConvLSTM, Shin et al. [25] used GNN along with Long Short-Term Memory (LSTM) and a data extraction approach, [26] which uses dynamic graph attention network (DGAT). However, from these other works, only Shin et al. [25] and Yu et al. [26] presented investigations on the NDT LCM, while none presented concept drift implementations.
Finally, in terms of the datasets used in these works, there is a noticeable concentration of studies [6, 22, 7, 25, 24] based on specific datasets derived from the OMNeT++ simulator [41]. This trend presents advantages and limitations. On the one hand, using datasets from a common source provides a stable baseline, enabling fair comparisons across different approaches in terms of generalization performance. On the other hand, relying solely on pre-built datasets limits the exploration of more realistic aspects of NDT, such as the mentioned closed-loop operations or digital twin synchronization, as there is no interaction with a real or emulated environment.
Given the current state of the literature on twin synchronization in the transport domain, Table 1 summarizes the works discussed in this section based on four primary characteristics: (i) closed-loop operation, referring to the interaction between the VTwin and PTwin; (ii) support for concept drift detection to handle traffic variability; (iii) the presence of twin synchronization mechanisms, regardless of whether they explicitly use concept drift detection; and (iv) consideration of NDT LCM elements, such as the planning, training, or operational phases. The comparative analysis in Table 1 emphasizes complementary and essential features of an NDT considered in this work, rather than focusing solely on improving the VTwin model itself, as is common in prior works. In this context, our objective is not to compare different VTwin models, but rather to consider a given VTwin model and examine how our proposed architecture aims to preserve its predictive performance through twin synchronization strategies
@@ 3 Enhanced Network Digital Twin
The proposed enhanced architecture is divided into the modules illustrated in Figure 1 and shares some inheritance from standardization bodies [34, 35] and industry players [2]. Following a bottom-up description, the first module corresponds to the PTwin defined by the switches in the transport network. Above the PTwin, we have the software-defined networking (SDN) controller and its related interfaces, north and southbound. Next to it, we have the VTwin, which in our solution is implemented using a GNN model. It provides a digital representation of the relationship between network traffic and topology links. This model plays a crucial role in estimating the total per-flow delay in the PTwin. The following module is the data management, which stores useful traffic database features (see Table 2). This module maintains a unified repository of transport network data, which is essential for feeding the VTwin with relevant features in the retraining process. Additionally, this module includes the necessary functions to collect and store traffic data, ensuring seamless integration with the overall system.
@@
Operating in the background to enable communication between these NDT modules, we have the interfaces defined by the SDN controller, which allow indirect data exchange between the VTwin and the controller, intended to apply any necessary modifications in the PTwin, and indirect data exchange between the database collector and the PTwin to fulfill this database with updated information from the physical environment. With this closed loop of control and traffic data between both twins, the VTwin predictions can be used in high-level applications, either with the involvement of a network operator, as defined by Aben-Athar et al. [4], or autonomously, without human supervision, as demonstrated by Messaoudi et al. [10]. In this article, we assume an SLA monitoring application. Complementary to this closed loop, the main contribution relies on the modules related to the synchronization of the NDT modules, which are responsible for synchronizing the status of the transport network of both domains in a reactive way. Along with the concept drift method, this module monitors if the incoming traffic that reaches the VTwin remains similar. Once changed and detected by the drift detector, this module has the capability to trigger a retraining process, since this change can also mean that the VTwin model is no longer predicting the total per-flow delay with an acceptable error performance. Finally, the source code of the proposed closed-loop architecture, its integration with concept drift techniques, the datasets used, and the evaluation methodology will be available in a public Git repository. [^1] https://github.com/lasseufpa/robust-ndt

### v2

@@ Abstract
The ability of the network digital twin (NDT) to remain aware of changes in its physical counterpart, known as the physical twin (PTwin), is a fundamental condition to enable timely synchronization, also referred to as twinning. In this way, considering a transport network, a key requirement is to handle unexpected traffic variability and dynamically adapt to maintain optimal performance in the associated virtual model, known as the virtual twin (VTwin). In this context, we propose a self-adaptive implementation of a novel NDT architecture designed to provide accurate delay predictions, even under fluctuating traffic conditions. This architecture addresses an essential challenge, underexplored in the literature: improving the resilience of data-driven NDT platforms against traffic variability and improving synchronization between the VTwin and its physical counterpart. Therefore, the contributions of this article rely on NDT lifecycle by focusing on the operational phase, where telemetry modules are used to monitor incoming traffic, and concept drift detection techniques guide retraining decisions aimed at updating and redeploying the VTwin when necessary. We validate our architecture with a network management use case, across various emulated network topologies, and diverse traffic patterns to demonstrate its effectiveness in preserving acceptable performance and predicting quality of service (QoS) metrics under unexpected traffic variation, such as delay and jitter. The results in all tested topologies, using the normalized mean square error as the evaluation metric, demonstrate that our proposed architecture, after a traffic concept drift, achieves a performance improvement in per-flow delay and jitter prediction of at least $64\%$ and $21\%$, respectively, compared to a configuration without NDT synchronization.
@@ 1 Introduction
NDT is a platform designed to aid in the management and automation of physical network communication tasks based on interactions with its virtual counterpart, also called virtual twin (VTwin), at a certain periodicity and fidelity [barricelli2019, ohlen2022]. This relation allows the network operator to emulate different scenarios in network planning to perform a what-if analysis [varela2023, abenathar2025]. This analysis aims to evaluate the impact of management changes in the network infrastructure when these tests are impossible due to logistical issues or costs that render them unfeasible, such as in production environments [Almasan2022DigitalTN, ohlen2022]. For that reason, this platform is considered an enabling technology to optimize 5th and 6th generation (5G/6G) network communication [ohlen2022], providing improved solutions to problems such as slicing management [Wang2022, farreras2025gnnetslice], routing optimization [rusek2020routenet, almasan2022], resource power allocation [varela2023], and admission control [messaoudi2023].
@@
Use of an NDT platform in a production environment can create several challenges, particularly concerning the long-term accuracy of the digital twin (DT). It is essential to account for variability in incoming data, ensuring that the VTwin continually adapts to reflect real-world changes. When we consider a VTwin neural network model to achieve this reliability, such as a graph neural network (GNN[ferriol2023routenet], it is fundamental to have a well-defined lifecycle management (LCM) that organizes all phases, including the planning, training, and operational aspects of the NDT platform [bariah2024interplay, Ahmadi2021, Kherbache2021]. When a deployed NDT is used during the operational phase, an important research question arises concerning the synchronization between the VTwin and the physical twin (PTwin). Synchronization issues are widely explored in the general context of DTs, considering applications at factory workstations, the industrial Internet of Things, power systems, etc. [Alghamdi2024, Hribernik2021]. However, these questions are rarely explored within the specific context of a transport network NDT based on GNN. Given that the transport network plays a crucial role as the interconnection between the core network and the radio access network (RAN[farreras2025gnnetslice, Morais2023], and the extensive body of literature supporting data-driven approaches using GNN [rusek2020routenet] for modeling and representing transport networks, it becomes possible to investigate particular challenges related to twin synchronization, such as the methods used to achieve it and the underlying synchronization patterns [Alghamdi2024].
In this sense, we propose an enhanced architecture grounded in the core lifecycle principles of NDT to address these challenges, focusing on maintenance and resilience. The architecture is designed to support a robust platform capable of adapting to traffic variability, also referred to as covariate drift, a special case of concept drift [baier2023, Ameur2025], which can arise from disruptions such as congestion, unreachable switches, broken links, or shifts in traffic patterns. Our approach emphasizes synchronization between the VTwin and the PTwin aimed to assist data-driven approaches remain up-to-date whenever a input distribution change is detected. We adopted a data-driven synchronization pattern [Alghamdi2024] managed through a concept drift detector that monitors network traffic for changes. When significant deviations are detected, a retraining process is triggered to realign the VTwin with the current state of the PTwin. Since our simulations employ a supervised learning setup, where the “ground truth” total delay and jitter are known, we evaluate and validate the generalization performance of the VTwin in the proposed architecture as the PTwin experiences dynamic changes in traffic intensity and patterns. The results of the proposed experiments demonstrate a performance improvement in all tested emulated topologies with different network traffic patterns, using the normalized mean square error (NMSE) as the evaluation metric. This shows that our proposed architecture achieves an improvement in total per-flow delay and jitter prediction of at least $64\%$ and $21\%$, respectively, after a traffic concept drift. In some other topologies, this improvement is greater than $99\%$ and $98\%$, respectively. Finally, with these improvements, we propose a final evaluation in a use case application based on service-level agreement (SLA) monitoring to demonstrate the importance of NDT synchronization, considering high-level metrics for network planning, such as the number of SLA violations.
@@
    A novel NDT architecture incorporating LCM principles during the operational phase, such as the re-deployment of an updated VTwin model when necessary. This contribution aimed to assist data-driven approaches (used to predict per-flow delay and jitter) remain up-to-date whenever a input distribution change is detected.
@@
    As part of our architecture, a concept drift detection module is integrated to monitor incoming network traffic features. This NDT module enables the analysis of traffic behavior to make informed decisions about when to update the VTwin model.
@@
    A use case application for network management related to SLA violation prediction compares the performance of the proposed application with and without NDT synchronization.
The remainder of this article is organized as follows: in Section 2 we compare our work with different proposed data-driven NDT in the literature, considering various levels of implementation in the transport network domain and their characteristics related to NDT supported modules, such as the synchronization module. In Section 3, we present the proposed enhanced NDT architecture, taking into account the main functional requirements of NDT, and show the essential building blocks of each, such as the VTwin architecture, the synchronization methods based on concept drift detection techniques, and complementary modules. In Section 4, we detail the experiments conducted to evaluate the proposed methods for twin synchronization and present the key evaluation results across various topologies and traffic patterns, as well as a use case with SLA monitoring. Finally, Section 5 summarizes the main contributions of this work and outlines directions for future research.
@@ 2 Related Work
Recent studies point DTs to an important technology for the 5G/6G networks [nguyen2021digital, khan2022digital, tao2024wireless, sheraz2024comprehensive, liu2024digital]. This aligns with their need to establish reliable environments for developing and testing their critical parts, such as wired environments [messaoudi2023], where DTs can provide excellent support. However, during this wave of adoption, the intersection of DT and networking—still an emerging area of study—led to a diverse range of approaches. Many of these did not align fully with the definitions of an NDT as systematized by standards organizations [itu_digital_twin, ndt_ietf] and industry stakeholders [ohlen2022, ORANRR]. Consequently, related work found in the literature presents limited analyzes of practical aspects. For example, regarding the system lifecycle or the twinning rates, the latter refers to the rate of synchronization between the PTwin current state and its corresponding representation in the VTwin [jones2020]. More specifically, much of the literature regarding NDTs of the transport domain has focused on using machine learning (ML)/surrogate model-based methods to generate the VTwin. The increase in popularity of this DT approach, which can also be called data-driven [bariah2024interplay], can be traced back to the work of Rusek et al. [rusek2020routenet], which established the usage of a GNN model called RouteNet for predicting path-level metrics in routing optimization and network upgrade use cases. The usage of this model was further investigated by several works, which were built on the previously established RouteNet [happ2021graph, saravanan2022enabling, Farreras2023, li2023learnable, ferriol2023routenet, messaoudi2023, Modesto2024].
Recently, the work of Ferriol-Galmés et al. [ferriol2023routenet] utilized a modified version of RouteNet, called RouteNet-Fermi, to model network communication with various topologies, traffic distributions, and queuing configurations, enabling the prediction of packet loss, delay, and jitter. In this work, since all data is provided from a packet-level simulator, there are no two-way interactions between PTwin and VTwin. Therefore, the case presented in the article does not constitute a closed-loop NDT. Moreover, the investigations did not address the possible performance degradation of the VTwin in cases of data and/or concept drifts [lu2018conceptdrift]. In this context, the cases presented in this work and several others built on RouteNet [happ2021graph, saravanan2022enabling, Farreras2023, li2023learnable, ferriol2023routenet] did not address the robustness aspect of an NDT, as they did not explore its lifecycle either.
The exception to this lack of closed-loop implementation in works based on RouteNet is the work of Messaoudi et al. [messaoudi2023], Zalat et al. [Zalat2024], and Aben-Athar et al. [abenathar2025], which are, to the best of our knowledge, the first closed-loop environments considering physical layers that interact with an artificial intelligence (AI) model. The works in [messaoudi2023, abenathar2025] feed an application module tailored to operate over the physical counterpart defined by virtual switches on Mininet [mininetsite]. The work in  [Zalat2024] took the same approach but considered a small-scale physical testbed. However, a noteworthy aspect of this work is that there are elements for twin synchronization that are based on a trigger-based online learning approach. The main difference between this work and our proposed architecture, apart from the use of online learning training, is that the authors specifically focused on exploiting the correlation between the predicted per-flow delay and the ground truth per-flow delay to update the states of the VTwin based on those of the PTwin. This correlation, when applied, can also be considered the simplest form to compare two twin states (outputs) without explicitly using a concept drift detector. However, a key drawback is the requirement for ground-truth labels to start a retraining process. For the other two works, despite closed-loop interactions between the physical and virtual domains, there is no DT lifecycle management controlling the rate at which both parts synchronize and update their respective states, for instance, through a retraining process. In this context, the authors in [messaoudi2023, abenathar2025] ran the traffic only once, and after completing it, they performed no further operations to detect whether the VTwin performance would remain acceptable.
In addition to these, other works in the literature either proposed purely GNN-based approaches, such as FlowDT [ferriol2022flowdt] or GNNetSlice [farreras2025gnnetslice], or adopted different ML models as their VTwins. For example, Lai et al. [lai2023deep] adopted a model named eConvLSTM; Shin et al. [shin2023network] used GNN along with Long Short-Term Memory (LSTM) and a data extraction approach [yu2023digital] that uses dynamic graph attention network (DGAT). However, among these other works, only Shin et al. [shin2023network] and Yu et al. [yu2023digital] presented investigations on the NDT LCM, while none presented concept drift implementations.
Finally, in terms of the datasets used in these works, there is a noticeable concentration of studies [Wang2022, saravanan2022enabling, farreras2025gnnetslice, shin2023network, Farreras2023] based on specific datasets derived from the OMNeT++ simulator [Varga2001]. This trend presents advantages and limitations. On the one hand, using datasets from a common source provides a stable baseline, enabling fair comparisons across different approaches in terms of generalization performance. On the other hand, relying solely on pre-built datasets limits the exploration of more realistic aspects of NDT, such as the aforementioned closed-loop operations or DT synchronization, as there is no interaction with a real or emulated environment.
Given the current state of the literature on twin synchronization in the transport network domain, Table 1 summarizes the works discussed in this section based on four primary characteristics: (i) closed-loop operation, referring to the interaction between the VTwin and PTwin; (ii) support for concept drift detection to handle traffic variability; (iii) the presence of twin synchronization mechanisms, regardless of whether they explicitly use concept drift detection; and (iv) consideration of NDT LCM elements, such as the planning, training, or operational phases. The comparative analysis in Table 1 emphasizes complementary and essential features of an NDT considered in this work, rather than focusing solely on improving the VTwin model itself, as is common in prior works. In this context, our objective is not to compare different VTwin models but rather to consider a given VTwin model and examine how our proposed architecture aims to preserve its predictive performance through a twin synchronization strategy. Finally, Table 1 considers only studies that explicitly use the term NDT. This decision reflects our intention to maintain terminological consistency. Although other platforms, such as cyber-physical system (CPS[kathiravelu2019sd, Wu2021], may share certain similarities with NDTs, they often involve broader specifications that extend beyond the scope of NDTs as assumed in our work.
@@
Table 1: Comparison between NDT characteristics of the related works and our proposed architecture with support for concept drift.
@@
Rusek et al., 2020 [rusek2020routenet]
Happ, et al., 2021 [happ2021graph]
Wang et al., 2022 [Wang2022]
Güemes-Palau et al., 2022 [guemes2022accelerating]
Ferriol-Galmés et al., 2022 [ferriol2022flowdt]
Saravanan et al., 2022 [saravanan2022enabling]
Lai et al., 2023 [lai2023deep]
Farreras et al., 2023 [Farreras2023]
Messaoudi et al., 2023 [messaoudi2023]
Shin et al., 2023 [shin2023network]
Yu et al., 2023 [yu2023digital]
Ferriol-Galmés et al., 2023 [ferriol2023routenet]
Zalat et al., 2024 [Zalat2024]
Modesto et al., 2024 [Modesto2024]
Farreras et al., 2025 [farreras2025gnnetslice]
Aben-Athar et al., 2025 [abenathar2025]
@@ 3 Enhanced Network Digital Twin
The proposed enhanced architecture is divided into the modules illustrated in Figure 1 and shares some inheritance from standardization bodies [itu_digital_twin, ndt_ietf] and industry players [ohlen2022]. Following a bottom-up description, the first module corresponds to the PTwin defined by the switches in the transport network. Above the PTwin, we have the software-defined networking (SDN) controller and its related interfaces, northbound and southbound. Next to it, we have the VTwin, which in our solution is implemented using a GNN model. It provides a digital representation of the relationship between network traffic and topology links. This model plays a crucial role in estimating the QoS parameters in the PTwin. Table 2 provides an overview of the terms used to represent both twins. The following module is the data management system, which stores useful features of the traffic database (see Table 3). This module maintains a unified repository of transport network data, which is essential for feeding the VTwin with relevant features in the retraining process. Additionally, this module includes the necessary functions to collect and store traffic data, ensuring seamless integration with the overall system.
Table 2: Description of terms used.
Term Description
$\mathcal{G}$ Transport network topology
$\mathcal{F}$ Set of network flows
$\mathcal{V}$ Set of nodes
$\mathcal{E}$ Set of links
$O$ Source switch
$D$ Destination switch
$e_{j}$ $j$-th link
$f_{t}$ $t$-th flow
$\mathbf{x}_{e_{j}}$ $j$-th link feature vector
$\mathbf{x}_{f_{t}}$ $t$-th flow feature vector
$x$ Average traffic rate
$T$ Total number of flows
$J$ Total number of links in a path
$P_{t}$ Probability distribution at discrete time $t$
$a(\cdot)$ Attention score
$\psi(\cdot)$ Message function
$\phi(\cdot)$ Update function
[…]
$\bigoplus$ Aggregation function
$\mathcal{O}$ Readout output
$\mathbf{j}$ Vector of link indices
$\mathbf{i}$ Vector of flow indices
$K$ Number of iteration on message passing
$\mathcal{E}^{\prime}$ Path of a specific flow
$p_{j}$ j-th link propagation delay
$q$ Type of QoS metric (delay or jitter)
$y^{(q)}$ per-flow QoS metric
$\hat{y}^{(q)}$ Predicted per-flow QoS metric
$m$ Flow embedding size
$n$ Link embedding size
$\mathbf{h}$ Embedding vector
$\mathcal{N}$ Neighborhood of a node
@@
Operating in the background to enable communication between these NDT modules, we have the interfaces defined by the SDN controller, which allow for indirect data exchange between the VTwin and the controller. This setup is intended to apply any necessary modifications in the PTwin and facilitate indirect data exchange between the database collector and the PTwin to fulfill this database with updated information from the physical environment. With this closed loop of control and traffic data between both twins, the VTwin predictions can be used in high-level applications, either with the involvement of a network operator, as defined by Aben-Athar et al. [abenathar2025], or autonomously, without human supervision, as demonstrated by Messaoudi et al. [messaoudi2023]. In this article, we assume an SLA monitoring application. Complementary to this closed loop, the main contribution relies on the modules related to the synchronization of the NDT modules, which are responsible for synchronizing the status of the transport network in both domains in a reactive manner. Along with the concept drift detector, this module monitors whether the incoming traffic that reaches the VTwin remains similar. Once changed and detected by the drift detector, this module has the capability to trigger a retraining process, as this change can also mean that the VTwin model is no longer predicting the total per-flow delay or jitter with acceptable error performance. Finally, the source code of the proposed closed-loop architecture, its integration with concept drift techniques, the datasets used, and the evaluation methodology are available in a public GitHub repository. [^1] https://github.com/lasseufpa/robust-ndt